# ChartEdit: How Far Are MLLMs From Automating Chart Analysis? Evaluating MLLMs’ Capability via Chart Editing

### v1

@@ Abstract
Although multimodal large language models (MLLMs) show promise in generating chart rendering code, chart editing presents a greater challenge. This difficulty stems from its nature as a labor-intensive task for humans that also demands MLLMs to integrate chart understanding, complex reasoning, and precise intent interpretation. While many MLLMs claim such editing capabilities, current assessments typically rely on limited case studies rather than robust evaluation methodologies, highlighting the urgent need for a comprehensive evaluation framework. In this work, we propose ChartEdit, a new high-quality benchmark designed for chart editing tasks. This benchmark comprises diverse editing instructions applied to real-world charts, with each instruction-chart instance having been manually annotated and validated for accuracy. Utilizing ChartEdit, we evaluate the performance of 10 mainstream MLLMs across two types of experiments, assessing them at both the code and chart levels. The results suggest that large-scale models can generate code to produce images that partially match the reference images. However, their ability to generate accurate edits according to the instructions remains limited. The state-of-the-art (SOTA) model achieves a score of only , highlighting significant challenges in precise modification. In contrast, small-scale models, including chart-domain models, struggle both with following editing instructions and generating overall chart images, underscoring the need for further development in this area. Code is available at https://github.com/xxlllz/ChartEdit.
@@
Xuanle Zhao1,111Equal contribution., Xuexin Liu1,111Equal contribution., Haoyue Yang1,111Equal contribution., Xianzhen Luo2, Fanhu Zeng1, Jianling Li 3, Qi Shi4222Corresponding author., Chi Chen4222Corresponding author. 1 Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 Harbin Institute of Technology, Harbin, China 3 Tianjin University, Tianjin, China 4 Tsinghua University, Beijing, China {zhaoxuanle2022, liuxuexin2022, yanghaoyue2024}@ia.ac.cn
@@ 1 Introduction
Recently, multimodal large language models (MLLMs) Liu et al. (2023a); Zhang et al. (2024a); Guo et al. (2025), which leverage the rich knowledge in large language models (LLMs) Dubey et al. (2024), have demonstrated impressive performance in processing and reasoning over various modalities, such as images Wang et al. (2024b); Zeng et al. (2024) and videos Li et al. (2024). However, much of the existing research in code generation has primarily focused on using text as the sole input Li et al. (2022); Luo et al. (2023), leaving the potential of multimodal information largely unexplored.
2 Related Works
2.1 Chart-Domain MLLMs
@@ 2.2 MLLMs For Code
Multimodal code generation has recently gained significant attention. Several works, like Design2Code Si et al. (2024) and Web2Code Yun et al. (2024), focus on assessing the capacities of MLLMs in generating HTML code for web pages. RoboCodeX Mu et al. (2024) proposes a multimodal code generation framework for robot behavior synthesis. The field of chart-to-code generation has also attracted considerable attention Zhao et al. (2025), as it focuses on generating code that accurately reproduces a given chart image. This task is challenging due to the complex visual elements in charts, demanding advanced models to accurately translate them into functional code. Recent works Wu et al. (2024); Xia et al. (2024); Zhang et al. (2024c) evaluate the capacities of MLLMs in this context. Although some works like Xia et al. (2024); Shi et al. (2024) have considered this problem, the editing instructions in their works only focus on one type of modification, which is not diverse enough.
@@ 3 ChartEdit
In this section, we first introduce the chart editing task and illustrate the data collection and generation pipeline.
@@ 3.3 Dataset Statistics and Analysis
Using LLM generation and manual annotation, we construct our evaluation benchmark, ChartEdit, which comprises <chart, code> pairs and <chart, instruction, code> triplets. We further analyze the diversity of ChartEdit from both chart and instruction perspectives. To ensure a diverse selection of chart types, we have predefined distinct chart categories, carefully curating instances for each type. The distribution of these chart types is provided in the Section A.2. Additionally, we enhance the diversity of editing instructions by classifying them into six categories: five generated by LLMs, as outlined in Section 3.2.3, and one written by humans. The proportions of each editing category are shown in Figure 3. In addition to exploring the diversity of editing types, we also examine the different formats in which these instructions are presented, including plain text descriptions, code symbols such as marker=’*’, and hexadecimal color codes like color=’#ABCDEF’. We embed these editing instructions and compare them with instructions from other chart-to-code benchmarks using dimensionality reduction techniques. The result in Figure 4 shows embedding after dimension reduction, which indicates that different editing types are generally separated, and the human-written instructions almost cover all the instruction types.
@@ 4.1 Baselines
For the evaluation, we use the direct instruction prompting method across all models. Detailed prompts and implementation steps are provided in the Figure 11.
@@ 4.2 Evaluation Metrics
Therefore, we propose evaluating both the code level and the chart level. The code-level evaluation is based on the source code, editing instructions, and output code. We use GPT-4o to score the models on two aspects: Modification Accuracy and Code Completeness. Building on related approaches Shi et al. (2024); Zhang et al. (2024c), we also evaluate the model’s overall generation capabilities at the chart level. This evaluation is performed by comparing how closely the generated chart reproduces the reference edited chart (manually created). The results reflect the degree of completeness of the generated chart image. Following Shi et al. (2024), we also utilize GPT-4o to evaluate six aspects: Types, Layout, Text, Data, Style, and Clarity. Detailed information is provided in the Figure 13.
@@ 4.3 Main Results
The main results for all the MLLMs are presented in Table 3. Although open-source models have made significant strides, and in some cases even outperformed GPT-4 in various tasks Masry et al. (2022); Zhang et al. (2024a), there is still a noticeable gap when it comes to handling complex multimodal tasks Wang et al. (2024c); Shi et al. (2024). In our experiments, GPT-4o delivers the best performance on the Chart w/ Code task, achieving the highest Execution Rate, Code-Level, and Chart-Level scores. However, the Code-Level score is still not high enough, indicating that GPT-4o faces challenges with precise editing. Among open-source MLLMs, InternVL2.5-78B achieves the best performance, but it still lags behind GPT in terms of code execution rate and chart-level metrics. Our analysis shows that the largest gap between proprietary and open-source models lies in generating the complete code corresponding to an image. There is little difference in instruction-following ability between the two, and both struggle with detailed modifications in chart editing. The chart-domain models perform worse in this task, and we believe this is due to the current limitations of MLLMs in the chart-domain, particularly in handling editing tasks. Although TinyChart Zhang et al. (2024b) and ChartLlama Han et al. (2023) have been fine-tuned on relevant datasets, both lack the ability to follow editing instructions effectively. The instructions in chart editing are much more diverse than those in direct chart-to-code generation tasks.
@@
The detailed scores of chart-level are presented in Table 4, and we select four models for analysis. Since the Chart w/o Code task is significantly more challenging, we only provide detailed scores for this task. The result shows that the performance gaps between proprietary and open-domain models are most noticeable in the text, data, and style aspects, which are key pieces of information in the chart. Table 5 presents the detailed scores at the code level, highlighting that there remains a gap between the state-of-the-art open-source models and proprietary models in generating complete code. At the same time, the large-scale model has a significant advantage over the small-scale model in terms of modification accuracy.
@@ 5 Discussion
In this section, we analyze the following questions.
RQ1: Is Chain-of-Thought (CoT) prompting useful for MLLMs in chart editing tasks for proprietary models? Answer: No. In the CoT setting, we provide MLLMs with zero-shot prompting, instructing them to first understand the source chart, then analyze the editing instructions, and finally generate the code. However, the result in Figure 5 shows whether using proprietary models or open-source models, CoT prompting shows almost no improvement in performance. We analyze the thought process of GPT-4o and find that while CoT helps the model better understand the components in the chart and the editing instructions, it does not result in more accurate code generation. The detailed CoT prompts are provided in the Figure 11.
RQ2: How does the performance of MLLMs vary across different types of editing instructions? In Figure 6, we compare the performance across different types of editing tasks and find that models are much more effective at generating code to modify chart types. Our analysis reveals that unlike tasks requiring first grounding and then modification of specific elements, the format conversion task focuses on altering the overall visual representation of the chart, which makes it easier to edit the code. However, at the chart level, model performance on format conversion tasks is limited. This suggests that while models can change the chart type effectively, they still struggle with more complex modifications involving other elements and fail to capture all the details of the chart.
RQ3: How does the performance of MLLMs vary across different types of charts? Figure 7 illustrates model performance across different chart types. The results show that the most significant performance discrepancy between open-source and proprietary models occurs with venn diagrams. Additionally, noticeable differences emerge in handling more complex chart types, such as errorbar and multi charts. Upon further analysis, we found that the comparable performance of InternVL2.5-7B and GPT-4o on the code-level tasks is due to the fact that this metric primarily evaluates the models’ capacity to follow instructions. Since most editing instructions for venn diagrams do not require altering the chart type itself, the performance gap at the code level is smaller than at the chart level.
RQ4: How do the error types of generated code vary across MLLMs? Figure 9 shows the types of runtime errors generated by different models. The results reveal that ValueError constitutes the largest proportion of errors, accounting for more than half of the total. Interestingly, the proportion of errors excluding ValueError tends to decrease as the model’s performance improves. Our analysis suggests that as code generation capabilities improve, many common errors are effectively mitigated while more complex issues related to data processing persist.
@@ A.1 Instruction Generation Details
In data generation, we use InternVL2.5-78B to evaluate whether the images crawled from Arxiv are charts and to score the quality of these chart images.
@@ A.2 Chart Statictic
We present the count of each chart type in Table 6. The largest proportions are observed in bar and line charts, which are the most common chart types in Arxiv papers.
@@
A.3 Prompts
A.4 Human Eval
We have evaluate four popular models on Chart w/o Code setting. This results are list below

### v2

@@ Abstract
Although multimodal large language models (MLLMs) show promise in generating chart rendering code, editing charts via code presents a greater challenge. This task demands MLLMs to integrate chart understanding and reasoning capacities, which are labor-intensive. While many MLLMs claim such editing capabilities, current evaluations rely on limited case studies, highlighting the urgent need for a comprehensive evaluation framework. In this work, we propose ChartEdit, a novel benchmark designed for chart editing tasks, featuring diverse editing instructions applied to real-world charts, each manually annotated and validated for accuracy. Utilizing ChartEdit, we evaluate the performance of 10 mainstream MLLMs across two types of experiments at both the code and chart levels. The results suggest that large-scale models can generate code to produce images that partially match the reference images. However, their ability to generate accurate edits according to the instructions remains limited. The state-of-the-art (SOTA) model achieves a score of only , highlighting significant challenges in precise modification. In contrast, small-scale models, including chart-domain models, struggle both with following editing instructions and generating overall chart images, underscoring the need for further development in this area. Code is available at https://github.com/xxlllz/ChartEdit.
@@
Xuanle Zhao1,111Equal contribution., Xuexin Liu1,111Equal contribution., Haoyue Yang1,111Equal contribution., Xianzhen Luo2, Fanhu Zeng1, Jianling Li3, Qi Shi4,222Corresponding authors., Chi Chen4,222Corresponding authors. 1 Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 Harbin Institute of Technology, Harbin, China 3 Tianjin University, Tianjin, China 4 Tsinghua University, Beijing, China {zhaoxuanle2022, liuxuexin2022, yanghaoyue2024}@ia.ac.cn
@@ 1 Introduction
Recently, multimodal large language models (MLLMs) Liu et al. (2023a); Zhang et al. (2024a); Guo et al. (2025); Zhang et al. (2025), which leverage the rich knowledge in large language models (LLMs) Dubey et al. (2024), have demonstrated impressive performance in processing and reasoning over various modalities, such as image Wang et al. (2024b); Yu et al. (2025), video Li et al. (2024) and speech Wang et al. (2024d). However, much of the existing research in code generation has primarily focused on using text as the sole input Li et al. (2022); Luo et al. (2023), leaving the potential of multimodal information largely unexplored.
2 Related Works
2.1 Chart-Domain MLLMs
@@ 2.2 MLLMs For Code
Multimodal code generation has recently gained significant attention. Several works, like Design2Code Si et al. (2024) and Web2Code Yun et al. (2024), focus on assessing the capacities of MLLMs in generating HTML code for web pages. RoboCodeX Mu et al. (2024) proposes a multimodal code generation framework for robot behaviour synthesis. The field of chart-to-code generation has also attracted considerable attention Zhao et al. (2025), as it focuses on generating code that accurately reproduces a given chart image. This task is challenging due to the complex visual elements in charts, demanding advanced models to accurately translate them into functional code. Recent works Wu et al. (2024); Xia et al. (2024); Zhang et al. (2024c) evaluate the capacities of MLLMs in this context. Although some works like Xia et al. (2024); Shi et al. (2024) have considered this problem, the editing instructions in their works only focus on one type of modification, which is not diverse enough.
@@ 3 ChartEdit
In this section, we first introduce the definition of the chart editing task and illustrate the data collection and organization pipeline of ChartEdit.
@@ 3.3 Dataset Statistics and Analysis
Using LLM generation and manual annotation, we construct our evaluation benchmark, ChartEdit, which comprises <chart, code> pairs and <chart, instruction, code> triplets. We further analyze the diversity of ChartEdit from both chart and instruction perspectives. To ensure a diverse selection of chart types, we have predefined distinct chart categories, carefully curating instances for each type. The distribution of these chart types is provided in the Section˜A.2. Additionally, we enhance the diversity of editing instructions by classifying them into six categories: five generated by LLMs, as outlined in Section˜3.2.3, and one written by humans. The proportions of each editing category are shown in Figure˜3. In addition to exploring the diversity of editing types, we also examine the different formats in which these instructions are presented, including plain text descriptions, code symbols such as marker=’*’, and hexadecimal color codes like color=’#ABCDEF’. We embed these editing instructions and compare them with instructions from other chart-to-code benchmarks using dimensionality reduction techniques. The result in Figure˜4 shows embedding after dimension reduction, which indicates that different editing types are generally separated, and the human-written instructions almost cover all the instruction types.
@@ 4.1 Baselines
For the evaluation, we use the direct instruction prompting method across all models. Detailed prompts and implementation steps are provided in the Figure˜11.
@@ 4.2 Evaluation Metrics
Therefore, we propose evaluating both the code level and the chart level. The code-level evaluation is based on the source code, editing instructions, and output code. We use GPT-4o to score the models on two aspects: Modification Accuracy and Code Completeness. Building on related approaches Shi et al. (2024); Zhang et al. (2024c), we also evaluate the model’s overall generation capabilities at the chart level. This evaluation is performed by comparing how closely the generated chart reproduces the reference edited chart (manually created). The results reflect the degree of completeness of the generated chart image. Following Shi et al. (2024), we also utilize GPT-4o to evaluate six aspects: Types, Layout, Text, Data, Style, and Clarity. Detailed information is provided in the Figure˜13.
@@ 4.3 Main Results
The main results for all the MLLMs are presented in Table˜3. Although open-source models have made significant strides, and in some cases even outperformed GPT-4 in various tasks Masry et al. (2022); Zhang et al. (2024a), there is still a noticeable gap when it comes to handling complex multimodal tasks Wang et al. (2024c); Shi et al. (2024). In our experiments, GPT-4o delivers the best performance on the Chart w/ Code task, achieving the highest Execution Rate, Code-Level, and Chart-Level scores. However, the Code-Level score is still not high enough, indicating that GPT-4o faces challenges with precise editing. Among open-source MLLMs, InternVL2.5-78B achieves the best performance, but it still lags behind GPT in terms of code execution rate and chart-level metrics. Our analysis shows that the largest gap between proprietary and open-source models lies in generating the complete code corresponding to an image. There is little difference in instruction-following ability between the two, and both struggle with detailed modifications in chart editing. The chart-domain models perform worse in this task, and we believe this is due to the current limitations of MLLMs in the chart-domain, particularly in handling editing tasks. Although TinyChart Zhang et al. (2024b) and ChartLlama Han et al. (2023) have been fine-tuned on relevant datasets, both lack the ability to follow editing instructions effectively. The instructions in chart editing are much more diverse than those in direct chart-to-code generation tasks.
@@
The detailed scores of the chart-level are presented in Table˜4, and we select four models for analysis. Since the Chart w/o Code task is significantly more challenging, we only provide detailed scores for this task. The result shows that the performance gaps between proprietary and open-domain models are most noticeable in the text, data, and style aspects, which are key pieces of information in the chart. Table˜5 presents the detailed scores at the code level, highlighting that there remains a gap between the state-of-the-art open-source models and proprietary models in generating complete code. At the same time, the large-scale model has a significant advantage over the small-scale model in terms of modification accuracy.
Furthermore, we also recruit human evaluators to score the editing results of four popular MLLMs. Details about the results are listed in Section˜A.3.
@@ 5 Discussion
In this section, we conduct many analyses to answer the following questions.
RQ1: Is Chain-of-Thought (CoT) prompting useful for MLLMs in chart editing tasks for proprietary models? Answer: No. In the CoT setting, we provide MLLMs with zero-shot prompting, instructing them to first understand the source chart, then analyze the editing instructions, and finally generate the code. However, the result in Figure˜5 shows whether using proprietary models or open-source models, CoT prompting shows almost no improvement in performance. We analyze the thought process of GPT-4o and find that while CoT helps the model better understand the components in the chart and the editing instructions, it does not result in more accurate code generation. The detailed CoT prompts are provided in the Figure˜11.
RQ2: How does the performance of MLLMs vary across different types of editing instructions? In Figure˜6, we compare the performance across different types of editing tasks and find that models are much more effective at generating code to modify chart types. Our analysis reveals that unlike tasks requiring first grounding and then modification of specific elements, the format conversion task focuses on altering the overall visual representation of the chart, which makes it easier to edit the code. However, at the chart level, model performance on format conversion tasks is limited. This suggests that while models can change the chart type effectively, they still struggle with more complex modifications involving other elements and fail to capture all the details of the chart.
RQ3: How does the performance of MLLMs vary across different types of charts? Figure˜7 illustrates model performance across different chart types. The results show that the most significant performance discrepancy between open-source and proprietary models occurs with venn diagrams. Additionally, noticeable differences emerge in handling more complex chart types, such as errorbar and multi charts. Upon further analysis, we found that the comparable performance of InternVL2.5-7B and GPT-4o on the code-level tasks is since this metric primarily evaluates the models’ capacity to follow instructions. Since most editing instructions for venn diagrams do not require altering the chart type itself, the performance gap at the code level is smaller than at the chart level.
RQ4: How do the error types of generated code vary across MLLMs? Figure˜9 shows the types of runtime errors generated by different models. The results reveal that ValueError constitutes the largest proportion of errors, accounting for more than half of the total. Interestingly, the proportion of errors excluding ValueError tends to decrease as the model’s performance improves. Our analysis suggests that as code generation capabilities improve, many common errors are effectively mitigated while more complex issues related to data processing persist.
@@ A.1 Instruction Generation Details
In data generation, we use InternVL2.5-78B to evaluate whether the images crawled from Arxiv are charts and to score the quality of these chart images. Charts with scores above 90 pass the initial filtering.
@@ A.2 Chart Statictic
We present the count of each chart type in Table˜6. The largest proportions are observed in bar and line charts, which are the most common chart types in Arxiv papers.
@@
A.3 Human Evaluation
We recruit human evaluators to evaluate four popular MLLMs on the Chart w/o Code setting with criteria denoted in Figure˜10. The results are listed below Table˜7, which show alignments between code-level with the human evaluations.
@@ A.3 Human Evaluation
A.4 Prompts
For the chart editing task, we utilize prompts in Figure˜11 to instruct MLLMs to generate the edited code. For evaluation metrics, we utilize prompts in Figure˜12 and Figure˜13 to evaluate the code-level and chart-level scores.